OpenStudy (anonymous):

simplify the expression \[9^{3/5} \over 9^2/5\]

OpenStudy (anonymous):

subtract the exponents

OpenStudy (anonymous):

\[\frac{9^{\frac{3}{5}}}{9^{\frac{2}{5}}}=9^{\frac{3}{5}-\frac{2}{5}}\]\]

OpenStudy (anonymous):

\[=9^{\frac{1}{5}}\]

OpenStudy (anonymous):

1 ?

OpenStudy (saifoo.khan):

Oh, when i saw the ques, the 2 seemed like 3. lol.

OpenStudy (saifoo.khan):

[MY BAD]/

OpenStudy (anonymous):

ahhhh